OpenStudy (anonymous):

algebraically solve the differential question of: dy/dx=(4+x^2)y

OpenStudy (anonymous):

this is a separable differential equation\[\frac{dy}{dx} = \left( 4 + x^2 \right)y\]\[\frac{dy}{y} = \left( 4 + x^2 \right)dx\]integrate both sides then its done